# Weyl semimetallic state with antiferromagnetic order in Rashba-Hubbard model

### v1

@@ I Introduction
In the two-dimensional systems, especially at the interface of heterostructures, absence of the inversion symmetry can generate the spin-orbit coupling (SOC), which is referred to as Rashba SOC rashba . The Rashba SOC depends linearly on the crystal momentum ${\bf k}$ and lifts the spin degeneracy of energy bands, and therefore can significantly influence the electronic and transport properties giving rise to a wide variety of fascinating properties bychkov ; rashba2 ; rashba3 . Moreover, by controlling the layer thickness of a hetro-structure or by the application of an external electric field, the strength of Rashba SOC can be controlled, making such systems more suitable for several potential technological applications nitta ; schapers ; bandyopadhyay ; jungwirth ; manchon ; chappert ; knill . This is unlike the centrosymmetric systems with transition metals as indispensable constituents, where the SOC results from the $LS$ coupling in the $d$ and $f$ orbitals.
@@
The SOC in a correlated-electron system introduces quantum frustration, which, in addition to affecting the nature of Mott transition, can stabilize plethora of exotic states of matter. These states of matter can include axion insulator, spin-orbit coupled Mott insulator, topological semimetals and insulators etc krempa ; dario ; manchon ; joel ; kubo . Although in non-centrosymmetric system, despite significant progress made in understanding different aspects of the Rashba SOC, relatively less attention has been paid to the consequences of its interplay with the electronic correlation. In a few steps taken in this direction, the nature of metal-to-insulator transition (MIT) and the phase diagram in the Rashba-Hubbard model were studied by using a variety of methods which include Hartree-Fock (HF) approximation kennedy , variational-Monte Carlo (VMC) kubo1 , cluster dynamical mean-field theory (CDMFT) xzhang etc.
@@
Sine-deformed mean-field theory (SDMFT) based on the HF approximation predicts an incommensurate spin-density wave (SDW) states even when the SOC may be vanishingly small and spiral orders for a larger SOC chisa_hota . Other HF-based theories suggest that a metallic state with antiferromagnetic (AFM) order for moderate electronic correlation, AFM insulator for strong electronic correlation, and a striped-magnetic order when both SOC and electronic correlations being large, are stabilized kennedy . The existence of AFM order for smaller SOC is also supported by CDMFT xzhang as well as plaquette-based investigations valentina . VMC studies point out the existence of Weyl semimetallic (WSM) state without AFM order for stronger SOC and electronic correlation kubo1 . Among various phases, thus, found, the metallic AFM state appears to be of considerable interest particularly given its pseuodgap like behavior. The current study attempts to provide a detailed investigation of this phase, most importantly, revealing its topological nature.
@@
Our finding suggests that for a reasonable range of SOC and on-site Coulomb interaction, a significant portion of the SOC vs interaction phase diagram is occupied by the metallic and insulating AFM ordered state with in-plane magnetic moment. The transition from the Rashba metal to AFM insulator does not happen directly upon increasing the electronic correlation parameter $U$. Instead, a WSM state with AFM order is stabilized in between the aforementioned two phases. The associated Weyl points (WPs) shift from the high-symmetry points $\Gamma$ and M in the paramagnetic Rashba metallic state to the points along the high-symmetry directions given by ±$\pm k_{x}\mp k_{y}=\pi$ and $k_{x}$ = -$k_{y}$ in the WSM-AFM state. Their respective winding numbers are 1 and -1. We also examine the spin-resolved quasiparticle intererence (QPI) patterns which is capable of revealing the nature of electronic state including the spin texture in the vicinity of the WPs.
@@ II.1 Model
[…]
@@ II.2 Hartree-Fock mean-field theory
${H}_{im}={-\frac{U}{2}\sum_{{\bf i}\sigma}\psi^{\dagger}_{\bf i}(\bf{\sigma}\cdot{\bf m}_{\bf i})\psi_{\bf i}},\vspace{-3mm}$ (7)
@@
Here, ${\bf\Delta}$ is the exchange field given by $2{\bf\Delta}=U{\bf m}$, where ${\bf m}=m_{x}\hat{x}+m_{y}\hat{y}+m_{z}\hat{z}$ represents the magnetic-moment vector in one of the sublattices. The direction of the magnetic-moment vector will be opposite in the other sublattice. The two-fold degenerate eigenvalues of the Hamiltonian [Eq. (8)] can be shown to be
$E_{k}$ $=$ ±[±$\pm\Bigg{[}\pm 2\bigg{(}\sqrt{\epsilon_{k}^{2}\beta_{k}^{2}+\epsilon_{k}^{2}\gamma_{k}^{2}+\beta_{k}^{2}\Delta_{y}^{2}+\gamma_{k}^{2}\Delta_{x}^{2}-2\beta_{k}\gamma_{k}\Delta_{x}\Delta_{y}}\bigg{)}$ (13)
@@ II.3 Magnetic susceptibility
In the previous subsection, we only discussed the Hamiltonian for the commensurate AFM order. Whether a commensurate or incommensurate AFM ordered state with in-plane or out-of-plane magnetic moments can be stabilized, it can be ascertained by examining the static susceptibility given by the 4$\times$4 matrix
@@ II.4 DOS modulation due to single impurity
Quasiparticle Interference (QPI) is a powerful technique to study the low-energy quasiparticle excitations in a system jennifer1 . The elastic scattering of quasiparticles by an impurity atom generates interference patterns, which may provide important insight especially into the electronic band structure in the vicinity of the Fermi level. The modulation in the local density of states (DOS) corresponding to the interference patterns can be calculated with the help of Green’s function.
@@
where $\hat{G}^{0}({\bf k},\omega)=(\hat{I}-\hat{\mathcal{H}}_{HF})^{-1}$ is the free particle Green’s function in AFM ordered state. $\hat{T}$ matrix of scattering by an impurity atom is defined in terms of the Green’s function given by
$\centering\hat{T}(\omega)=(\hat{{I}}-\mathcal{\hat{G}}^{0}(\omega))^{-1}\hat{V}^{imp},\@add@centering$ (20)
@@ III.1 Magnetic-ordering instability
We begin by examining the magnetic ordering instability in the Rashba-Hubbard model. Fig. 1 shows the static magnetic susceptibilities in different directions for the unordered state. Because of the SOC, $SU(2)$ rotational symmetry is absent, therefore, the in-plane component $\chi^{0}_{xx}(\bf q)$ and the out-of-plane component $\chi^{0}_{zz}(\bf q)$ are different. Moreover, $\chi^{0}_{xx}(\bf q)$ and $\chi^{0}_{yy}(\bf q)$ are also different along the high-symmetry directions such as $\Gamma$-X, $\Gamma$-Y, M-Y etc. We find that both $\chi^{0}_{xx}(\bf q)$ or $\chi^{0}_{yy}(\bf q)$ show stronger divergences at ($\pi,\pi$) as compared to $\chi^{0}_{zz}(\bf q)$, when $\lambda$ is small, implying that the magnetic moments will preferably be oriented in the $x-y$ plane. Secondly, upon increasing $\lambda$, peak position for $\chi^{0}_{zz}(\bf q)$ shifts away from $(\pi,\pi)$ to an incommensurate wavevector. There is no similar shift in the peak position for the in-plane component of the magnetic susceptibility. The latter is a robust feature in a realistic range for the SOC parameter, which is a result of nesting between the Fermi pockets around $\Gamma$ and M points with nesting vector ($\pi,\pi$) [Fig. 2]. On the other hand, shifting of $\chi^{0}_{zz}(\bf q)$ peak indicates that this component of susceptibility is likely to arise from the intrapocket nesting.
[…]
Figure 2: Fermi-surfaces with one of the prominent nesting vectors ${\bf Q}=(\pi,\pi)$ at half-filling for (a) $\lambda=0.4$ and (b) $\lambda=0.8$, where the range of both $k_{x}$ and $k_{y}$ is $[-\pi,\pi]$
@@ III.2 WSM-AFM state
Having examined the robustness of the AFM state with in-plane magnetic moment, we now explore the phase diagram in the $U-\lambda$ parameter space, by using the self-consistent mean-field theoretic approach discussed above. Since the in-plane RPA-level magnetic susceptibility is expected to show divergence at the wavevector ($\pi,\pi$), the initial value of out-of-plane magnetic moment, i.e., $m_{z}$ is set to be zero without any loss of generality. Only two types of phases are expected to occur, one with self-consistently obtained magnetic order, and another without magnetic order. Throughout the calculations, the inverse temperature parameter ($\beta=1/kT)$ is fixed to be 1000 with $k$ as the Boltzmann constant, which corresponds to $\sim 5$K, if tsimilar-$t\sim$ 0.5eV.
@@
All the terms in the Hamiltonian do not commute even for $\Delta_{x}=\Delta_{y}=0$, i. e., when the AFM order is absent. In that case, along ±$\pm k_{x}\mp k_{y}=\pi$, the first term contribution vanishes, the remaining two terms anticommute. Then, the Hamiltonian has both time-reversal ($\mathcal{T}=i\sigma_{2}K$) and inversion ($\mathcal{P}=\tau_{1}$) symmetries, which leads to the two-fold degenerate bands along (0, $\pi$) $\rightarrow$ (-$\pi$/2, $\pi/2$) and Dirac points (DPs) at (0, $\pi$) and ($\pi$, 0). This is mainly the consequence of Brillouin zone folding. In the absence of magnetic ordering, the folding is not required, then the degeneracy disappears and the DPs change into the WPs. The degerancy of the bands is lifted and the DPs are split into WPs if the system develops non-zero magnetic moment. One of the WPs is shifted away from (0, $\pi$) along $(0,\pi)\rightarrow(-\pi/2,\pi/2)$. Another band degeneracy occurs along ($\pi,0$) $\rightarrow$ (0, $\pi$), i.e., along $k_{x}+k_{y}=\pi$, which persists even in the AFM ordered state. This is a consequence of the fact that the first term vanishes while the rest anticommute with each other.
@@
The bands crossing at the Fermi level in the WSM-AFM state disappears with a rise in the value of $U$. The boundary between the AFM-I and WSM-AFM state is determined by the condition $|\Delta|\leq 2\lambda$, which should be satisfied by the magnetic exchange coupling $\Delta$. On lowering $U$ further, there is a phase transition from WSM-AFM to RM with the magnetic moment vanishing to zero in the self-consistent scheme. In the following, we obtain the condition to be satisfied by the magnetic order parameter to determine if the AFM ordered state is Weyl semimetallic or not.
@@
First of all, we note that the self-consistently obtained in-plane magnetization has the same magnitude along both directions, i.e., $\Delta_{x}=\Delta_{y}=\Delta$. If the condition $|\Delta|\leq 2\lambda$ is satisfied then one pair of WPs ($W_{1}$) is located along $k_{x}-k_{y}=\pi$ and $-k_{x}+k_{y}=\pi$ directions whereas the other pair ($W_{2}$) along $k_{x}=-k_{y}$ as seen in Fig. 5. Note that the word “pair” does not necessarily mean that they have opposite winding number as to be discussed later, it is used here only to indicate their location in the Brillouin zone.
III.3 WPs along ±$\pm k_{x}\mp k_{y}=\pi$
@@
For the pair of Weyl points $W_{1}$ located along $k_{x}-k_{y}=\pi$ and $-k_{x}+k_{y}=\pi$ [Fig. 5], $\epsilon_{\bf k}=0$ and $\beta_{\bf k}=-\gamma_{\bf k}$. Therefore, the eigenvalues [Eq. (9)] of the Hamiltonian matrix in the magnetically ordered state reduce to
@@ III.5 Chern numbers for WPs $W_{1}$ and $W_{2}$
$|u_{n}({\bf k})\rangle$ is the eigenvector of the Hamiltonian given by Eq. (8). The calculation by using Eq. (29) yields the winding number $w$ for WPs $W_{1}$ and $W_{2}$ along ±$\pm k_{x}\mp k_{y}=\pi$ and $k_{x}=-k_{y}$, 1 and -1, respectively. Fig. 7 shows the vector field associated with the Berry connection in the entire Brillouin zone while a zoomed in view for two such points with anticlockwise and clockwise rotations is presented in Fig. 8.

### v2

@@ I Introduction
In the two-dimensional systems, especially at the interface of heterostructures, absence of the inversion symmetry can generate the spin-orbit coupling (SOC), which is referred to as Rashba SOC rashba . The Rashba SOC depends linearly on the crystal momentum $\bf k$} and lifts the spin degeneracy of energy bands, and therefore can significantly influence the electronic and transport properties giving rise to a wide variety of fascinating properties bychkov ; rashba2 ; rashba3 . Moreover, by controlling the layer thickness of the heterostructure or by the application of an external electric field, the strength of Rashba SOC can be controlled, making such systems more suitable for several potential technological applications nitta ; schapers ; bandyopadhyay ; jungwirth ; manchon ; chappert ; knill . This is unlike the centrosymmetric systems with transition metals as indispensable constituents, where the SOC results from the $LS$ coupling in the $d$ and $f$ orbitals.
@@
The SOC in a correlated-electron system introduces quantum frustration, which, in addition to affecting the nature of Mott transition, can stabilize plethora of exotic states of matter. These states of matter can include axion insulator, spin-orbit coupled Mott insulator, topological semimetals and insulators etc krempa ; dario ; jacob ; manchon ; joel ; kubo . Although in non-centrosymmetric system, despite significant progress made in understanding different aspects of the Rashba SOC, relatively less attention has been paid to the consequences of its interplay with the electronic correlation. In a few steps taken in this direction, the nature of metal-to-insulator transition (MIT) and the phase diagram in the Rashba-Hubbard model were studied by using a variety of methods which include Hartree-Fock (HF) approximation kennedy , variational-Monte Carlo (VMC) kubo1 , cluster dynamical mean-field theory (CDMFT) xzhang etc.
@@
Sine-deformed mean-field theory (SDMFT) based on the HF approximation predicts an incommensurate spin-density wave (SDW) states even when the SOC may be vanishingly small and spiral orders for a larger SOC chisa_hota . Other HF-based theories suggest that a metallic state with antiferromagnetic (AFM) order for moderate electronic correlation, AFM insulator for strong electronic correlation, and a striped-magnetic order when both SOC and electronic correlations being large, are stabilized kennedy . The existence of AFM order for smaller SOC is also supported by CDMFT xzhang as well as plaquette-based investigations valentina . VMC studies point out the existence of Weyl semimetallic (WSM) state without AFM order for stronger SOC and electronic correlation kubo1 . Among various phases, thus, found, the metallic AFM state appears to be of considerable interest particularly given its pseudogap like behavior. The current study attempts to provide a detailed investigation of this phase, most importantly, revealing its topological nature.
@@
Our finding suggests that for a reasonable range of SOC and on-site Coulomb interaction, a significant portion of the SOC vs interaction phase diagram is occupied by the metallic and insulating AFM ordered state with in-plane magnetic moment. The transition from the Rashba metal to AFM insulator does not happen directly upon increasing the electronic correlation parameter $U$. Instead, a WSM state with AFM order is stabilized in between the aforementioned two phases. The associated Weyl points (WPs) shift from the high-symmetry points $\Gamma$ and M in the paramagnetic Rashba metallic state to the points along the high-symmetry directions given by $\pm k_{x}\mp k_{y}=\pi$ and $k_{x}$ = -$k_{y}$ in the WSM-AFM state. Their respective winding numbers are 1 and -1. We also examine the spin-resolved quasiparticle interference (QPI) patterns which is capable of revealing the nature of electronic state including the spin texture in the vicinity of the WPs.
@@ II.1 Model
$H_{R}=\lambda\sum_{{\bf i},\sigma,\sigma^{\prime}}[\textit{i}(d^{\dagger}_{{\bf i},\sigma}\sigma_{\sigma\sigma^{\prime}}^{x}d_{{\bf i}+\hat{y},\sigma^{\prime}}\\ -d^{\dagger}_{{\bf i},\sigma}\sigma_{\sigma\sigma^{\prime}}^{y}d_{{\bf i}+\hat{x},\sigma^{\prime}})+h.c.],$ (4)
@@ II.2 Hartree-Fock mean-field theory
${H}_{im}={-\frac{U}{2}\sum_{{\bf i}\sigma}\psi^{\dagger}_{\bf i}(\bf{\sigma}\cdot{\bf m}_{\bf i})\psi_{\bf i}},\vskip-8.53581pt$ (7)
@@
Here, $\bf\Delta$} is the exchange field given by $2{\bf\Delta}=U{\bf m}$, where $^{\bf m}=m_{x}\hat{x}+m_{y}\hat{y}+m_{z}\hat{z}$ represents the magnetic-moment vector in one of the sublattices. The direction of the magnetic-moment vector will be opposite in the other sublattice. The two-fold degenerate eigenvalues of the Hamiltonian [Eq. (8)] can be shown to be
$E_{k}$ $=$ $\pm\Bigg[\pm 2\bigg(\sqrt{\epsilon_{k}^{2}\beta_{k}^{2}+\epsilon_{k}^{2}\gamma_{k}^{2}+\beta_{k}^{2}\Delta_{y}^{2}+\gamma_{k}^{2}\Delta_{x}^{2}-2\beta_{k}\gamma_{k}\Delta_{x}\Delta_{y}}\bigg)$ (9)
@@ II.3 Magnetic susceptibility
In the previous subsection, we only discussed the Hamiltonian for the commensurate AFM order. Whether a commensurate or incommensurate AFM ordered state with in-plane or out-of-plane magnetic moments can be stabilized, it can be ascertained by examining the static susceptibility given by the $\times4$ matrix
@@ II.4 DOS modulation due to single impurity
Quasiparticle interference is a powerful technique to study the low-energy quasiparticle excitations in a system jennifer1 . The elastic scattering of quasiparticles by an impurity atom generates interference patterns, which may provide important insight especially into the electronic band structure in the vicinity of the Fermi level. The modulation in the local density of states (DOS) corresponding to the interference patterns can be calculated with the help of Green’s function.
@@
where $\hat{G}^{0}({\bf k},\omega)=(\omega\hat{I}-\hat{\mathcal{H}}_{HF})^{-1}$ is the free particle Green’s function in AFM ordered state. $\hat{T}$ matrix of scattering by an impurity atom is defined in terms of the Green’s function given by
$\centering\hat{T}(\omega)=(\hat{{I}}-\hat{V}^{imp}\mathcal{\hat{G}}^{0}(\omega))^{-1}\hat{V}^{imp},\@add@centering$ (16)
@@ III.1 Magnetic-ordering instability
We begin by examining the magnetic ordering instability in the Rashba-Hubbard model. Fig. 1 shows the static magnetic susceptibilities in different directions for the unordered state. Because of the SOC, $SU(2)$ rotational symmetry is absent, therefore, the in-plane component $\chi^{0}_{xx}(\bf q)$ and the out-of-plane component $\chi^{0}_{zz}(\bf q)$ are different. Moreover, $\chi^{0}_{xx}(\bf q)$ and $\chi^{0}_{yy}(\bf q)$ are also different along the high-symmetry directions such as $\Gamma$-X, $\Gamma$-Y, M-Y etc. We find that both $\chi^{0}_{xx}(\bf q)$ or $\chi^{0}_{yy}(\bf q)$ show stronger divergences at ($\pi,\pi$) as compared to $\chi^{0}_{zz}(\bf q)$, when $\lambda$ is small, implying that the magnetic moments will preferably be oriented in the xyx-y plane. Secondly, upon increasing $\lambda$, peak position for $\chi^{0}_{zz}(\bf q)$ shifts away from $(\pi,\pi)$ to an incommensurate wavevector. There is no similar shift in the peak position for the in-plane component of the magnetic susceptibility. The latter is a robust feature in a realistic range for the SOC parameter, which is a result of nesting between the Fermi pockets around $\Gamma$ and M points with nesting vector ($\pi,\pi$) [Fig. 2]. In the large SOC limit, $t/\lambda\rightarrow$ 0, we find that $\chi_{xx}$ and $\chi_{yy}$ are peaked at $(0,\pi)$ and $(\pi,0)$, respectively, whereas $\chi_{zz}$ at (0, 0) instead. Interestingly, all the three peaks have the same height indicating the same strength of instability against ferromagnetic or stripe order. The former set of peaks arise from the interpocket scattering between the pockets around $(0,0)$ and ($\pi,\pi$). The latter peak originates from the intra pocket scattering. It may be noted that, in the $t/\lambda\rightarrow 0$, the tiny pockets formed around $(0,0)$ and ($\pi,\pi$) becomes Weyl points kubo2 .
Figure 1: Various components of the static spin susceptibility in the whole Brillouin zone $(a$-$f)$ and also along the high-symmetry directions $(g$-$h)$ showing peaks for $\lambda=0.3$ and $0.6$.
Figure 2: Fermi-surfaces with one of the prominent nesting vectors ${\bf Q}=(\pi,\pi)$ at half-filling for (a) $\lambda=0.4$ and (b) $\lambda=0.8$, where the range of both $k_{x}$ and $k_{y}$ is $\pi,\pi]$ .
@@ III.2 WSM-AFM state
Having examined the robustness of the AFM state with in-plane magnetic moment, we now explore the phase diagram in the $U-\lambda$ parameter space, by using the self-consistent mean-field theoretic approach discussed above. Since the in-plane RPA-level magnetic susceptibility is expected to show divergence at the wavevector ($\pi,\pi$), the initial value of out-of-plane magnetic moment, i.e., $m_{z}$ is set to be zero without any loss of generality. Only two types of phases are expected to occur, one with self-consistently obtained magnetic order, and another without magnetic order. Throughout the calculations, the inverse temperature parameter ($\beta=1/kT)$ is fixed to be 1000 with $k$ as the Boltzmann constant, which corresponds to $\sim 5$K, if $t\sim$ 0.5eV.
@@
All the terms in the Hamiltonian do not commute even for $\Delta_{x}=\Delta_{y}=0$, i. e., when the AFM order is absent. In that case, along $\pm k_{x}\mp k_{y}=\pi$, the first term contribution vanishes, the remaining two terms anticommute. Then, the Hamiltonian has both time-reversal ($\mathcal{T}=i\sigma_{2}K$) and inversion ($\mathcal{P}=\tau_{1}$) symmetries, which leads to the two-fold degenerate bands along (0, $\pi$) $\rightarrow$ (-$\pi$/2, $\pi/2$) and Dirac points (DPs) at (0, $\pi$) and ($\pi$, 0). This is mainly the consequence of Brillouin zone folding. In the absence of magnetic ordering, the folding is not required, then the degeneracy disappears and the DPs change into the WPs. The degerancy of the bands is lifted and the DPs are split into WPs if the system develops non-zero magnetic moment. One of the WPs is shifted away from (0, $\pi$) along $(0,\pi)\rightarrow(-\pi/2,\pi/2)$. Another band degeneracy occurs along ($\pi,0$) $\rightarrow$ (0, $\pi$), i.e., along $k_{x}+k_{y}=\pi$, which persists even in the AFM ordered state. This is a consequence of the fact that the first term vanishes while the rest anticommute with each other.
@@
The bands crossing at the Fermi level in the WSM-AFM state disappears with a rise in the value of $U$. The boundary between the AFM-I and WSM-AFM state is determined by the condition $|\Delta|\leq 2\lambda$, which should be satisfied by the magnetic exchange coupling $\Delta$. On lowering $U$ further, there is a phase transition from WSM-AFM to RM with the magnetic moment vanishing to zero in the self-consistent scheme. This happens because the spin susceptbility does not diverge at $\bf Q$} for a given set of parameters, which includes on-site Coulomb interaction. In the following, we obtain the condition to be satisfied by the magnetic order parameter to determine if the AFM ordered state is Weyl semimetallic or not.
@@
First of all, we note that the self-consistently obtained in-plane magnetization has the same magnitude along both directions, i.e., $\Delta_{x}=\Delta_{y}=\Delta$. If the condition $|\Delta|\leq 2\lambda$ is satisfied then one pair of WPs ($W_{1}$) is located along $k_{x}-k_{y}=\pi$ and $k_{x}+k_{y}=\pi$ directions whereas the other pair ($W_{2}$) along $k_{x}=-k_{y}$ as seen in Fig. 5. Note that the word “pair” does not necessarily mean that they have opposite winding number as to be discussed later, it is used here only to indicate their location in the Brillouin zone.
III.3 WPs along $\pm k_{x}\mp k_{y}=\pi$
@@
For the pair of Weyl points $W_{1}$ located along $k_{x}-k_{y}=\pi$ and $k_{x}+k_{y}=\pi$ [Fig. 5], $\epsilon_{\bf k}=0$ and $\beta_{\bf k}=-\gamma_{\bf k}$. Therefore, the eigenvalues [Eq. (9)] of the Hamiltonian matrix in the magnetically ordered state reduce to
@@ III.5 Chern numbers for WPs $W_{1}$ and $W_{2}$
$|u_{n}({\bf k})\rangle$ is the eigenvector of the Hamiltonian given by Eq. (8). The calculation by using Eq. (29) yields the winding number $w$ for WPs $W_{1}$ and $W_{2}$ along $\pm k_{x}\mp k_{y}=\pi$ and $k_{x}=-k_{y}$, 1 and -1, respectively. Fig. 7 shows the vector field associated with the Berry connection in the entire Brillouin zone while a zoomed in view for two such points with anticlockwise and clockwise rotations is presented in Fig. 8.